# T2IShield: Defending Against Backdoors on Text-to-Image Diffusion Models

### v1

@@ Abstract
However, very recently proposed methods [44, 25, 8, 48, 52] show that T2I diffusion models are vulnerable to backdoor attacks. The attacker aims to manipulate the infected T2I diffusion model to generate a specified content caused by a pre-defined word (trigger), while maintaining good performance on benign inputs. The infected model can be maliciously used to generate taboo content and illegal watermarks. With the increasing number of pretrained T2I diffusion models downloaded from open-source websites like Civitai [1] by users and institutions, it becomes crucial to tell if these models suffered from a backdoor attack.
@@ 3 Methods
In this section, we introduce the details of our T2IShield, which is a comprehensive defense method to detect, localize, and mitigate backdoor attacks on T2I diffusion models. We firstly give a brief overview of our method. Then, we discuss cross-attention maps of the T2I diffusion model, which plays a key role in the detection algorithm. Finally, we introduce defense methods for detection, localization, and mitigation.
@@ 3.1 Overview of Our Methods
Detection: Given a T2I diffusion model downloaded from the untrustworthy third-party platforms and a set of input prompts to be tested, we aim to detect backdoor samples. We find that the backdoor trigger assimilates the cross-attention maps of other tokens, which we refer to as the "Assimilation Phenomenon". By modeling the structural correlation of the attention maps, T2IShield conducts two real-time binary classification methods, i.e., Frobenius Norm Threshold Truncation (FTT) and Covariance Discriminant Analysis (CDA),
@@ 3.2 Assimilation Phenomenon
Key Intuition. We visualize the attention maps of a benign sample and two backdoor samples [44, 8] in Fig. 1. The key observation is that the backdoor trigger will assimilate other tokens. Intuitively, in order to generate a specific content, the trigger must suppress the representation of other tokens. On the contrary, for a benign sample, each token’s attention map strongly correlates with its semantic information [21], e.g., pixels of the glasses are correlated with the word "glasses". By leveraging the difference of the structural correlation between attention maps, we propose two simple but effective backdoor detection methods, i.e., F Norm Threshold Truncation (FTT) and Covariance Discriminative Analysis (CDA).
@@ 3.3 Backdoor Detection
Covariance Discriminative Analysis. Since F norm only reflects the dispersion of the set of attention maps , leading to a coarse-grained representation of the correlation within the data. We then conduct classification by computing their covariance as a fine-grained feature. Inspired by image set classification [50, 26], we regard the set of attention maps as an image set and treat the problem as classifying points on a Riemannian manifold formed by Symmetric Positive Definite (SPD) matrices.
Formally, given a sequence of attention maps , we first flatten each attention map to , . Considering that attention maps are very sparse, we perform PCA (Principal Component Analysis) to reduce the dimension for each set of attention maps first and get , . Here, is the hyperparameter for the principal dimension.
@@ 4.1 Settings
Attack Models. We consider two types of attack models in the experiment, where Rickrolling [44] leverages the vulnerability of the text encoder (i.e., CLIP [38]) and Villan Diffusion [8] leverages the vulnerability of the UNet [41]. Followed the original settings [44, 8], we use stable diffusion v1.4 [39] as the T2I pre-trained model. For each backdoor attack method, we train eight types of backdoors. In particular, in Rickrolling, we set the loss weight to and fine-tune the encoder for 100 epochs with a clean batch size of 64. In Villan Diffusion, we fine-tune the model on CelebA-HQ-Dialog dataset [27] with LoRA [24] rank as 4 and the training batch size as 1. All experiments are conducted on RTX 4090 GPU with 24 GB memory.
@@
Evaluation Settings. Detection: We select 3000 backdoor samples which contains eight types of triggers and 3000 benign samples for training. The test dataset consists of 3000 backdoor samples which contains other eight types of triggers and 3000 other benign samples. Besides, the target content of the triggers are also different between the training and test datasets. We compare each backdoor detection method in terms of precision, recall, and F1 score, respectively. We also report the inference time for each method. Localization: We test each backdoor attack method with eight different triggers. We conduct experiments to analyze the impact of different similarity thresholds and different similarity calculation tools (i.e., CLIP [38] and DinoV2 [37]). A total of 1000 backdoor samples and 1000 benign samples are used. F1 score is employed as the evaluation metric for each localization method. The trigger length of Algorithm 1 is set to 1 as we primarily focus on characters-lever trigger [9]. Mitigation: We test each backdoor method with eight different triggers, and each trigger includes 100 backdoor samples. We utilize two state-of-the-art concept editing methods for mitigating backdoors. The Attack Success Rate (ASR) is computed on each trigger, which is the proportion of the prompt containing the trigger that generates the content specified by the attacker. Besides, we design a novel metric called Average Similarity to Benign (ASB), where we utilize CLIP [38] to compute the image similarity between the benign outputs and the mitigated outputs. We consider ASB is a more strict metric to evaluate the effectiveness of the backdoor mitigation on T2I diffusion models.
@@ 4.2 Detection Results
Results. Tab. 1 and Tab. 2 show the detection performances of F Norm Threshold Truncation (FTT) and Covariance Discriminative Analysis (CDA) under two attack scenarios, respectively. We find that backdoor samples from Villan Diffusion [8] are more deceptive than samples from Rickrolling [44]. This is reflected in the lower F1 score obtained by both detection methods when applied to samples from Villan Diffusion. As shown in Fig. 1, triggers from Rickrolling simultaneously assimilates the structure and intensity of attention maps from other tokens. In contrast, although triggers from Villan Diffusion also exhibit an "Assimilation Phenomenon", each token has variations in the response intensity. This leads to a closer resemblance to the response patterns of benign samples. Besides, CDA shows more robust detection performance on two backdoor attack scenarios, where it all achieves an average F1 score of 88.9 detection results compared to an average F1 score of 86.5 by FTT. In particular, for detecting hard samples from Villan Diffusion, it shows an improvement of 8.9 F1 score compared to the FTT method.
Besides, we record the average inference time of the FTT and CDA for detecting a sample on RTX 4090 32GB GPU. As can be seen, both of them perform in real-time, i.e., 9.4 ms for FTT and 11.7 ms for CDA, achieving low computational cost.

### v2

@@ Abstract
However, very recently proposed methods [43, 24, 8, 47, 51] show that T2I diffusion models are vulnerable to backdoor attacks. The attacker aims to manipulate the infected T2I diffusion model to generate a specified content caused by a pre-defined word (trigger), while maintaining good performance on benign inputs. The infected model can be maliciously used to generate taboo content and illegal watermarks. With the increasing number of pretrained T2I diffusion models downloaded from open-source websites [1] by users and institutions, it becomes crucial to tell if these models suffered from a backdoor attack.
@@ 3 Methods
In this section, we introduce the details of our T2IShield. We firstly give a brief overview of our method. Then, we discuss cross-attention maps of the T2I diffusion model, which plays a key role in the detection algorithm. Finally, we introduce defense methods for detection, localization, and mitigation.
@@ 3.1 Overview of Our Methods
Detection: Given a T2I diffusion model downloaded from the untrustworthy third-party platforms and a set of input prompts to be tested, we aim to detect backdoor samples. We find that the backdoor trigger assimilates the cross-attention maps of other tokens, which we refer to as the "Assimilation Phenomenon". By modeling the structural correlation of the attention maps, T2IShield conducts two real-time binary classification methods, i.e., Frobenius Norm Threshold Truncation (FTT) and Covariance Discriminant Analysis (CDA).
@@ 3.2 Assimilation Phenomenon
Key Intuition. We visualize the attention maps of a benign sample and two backdoor samples [43, 8] in Fig. 1. The key observation is that the backdoor trigger assimilates other tokens. Intuitively, in order to generate a specific content, the trigger must suppress the representation of other tokens. On the contrary, for a benign sample, each token’s attention map strongly correlates with its semantic information [20]. By leveraging the difference of the structural correlation between attention maps, we propose two simple but effective backdoor detection methods, i.e., F Norm Threshold Truncation and Covariance Discriminative Analysis.
@@ 3.3 Backdoor Detection
Covariance Discriminative Analysis. Since F norm only reflects the dispersion of the set of attention maps , leading to a coarse-grained representation of the correlation within the data. We then conduct classification by computing their covariance as a fine-grained feature. Inspired by image set classification [49, 25], we regard the set of attention maps as an image set and treat the problem as classifying points formed by Symmetric Positive Definite matrices.
@@ 4.1 Settings
Attack Models. We consider two types of attack models in the experiment, where Rickrolling [43] leverages the vulnerability of the text encoder (i.e., CLIP [37]) and Villan Diffusion [8] leverages the vulnerability of the UNet [40]. Followed the original settings [43, 8], we use stable diffusion v1.4 [38] as the T2I pre-trained model. For each backdoor attack method, we train eight types of backdoors. In particular, in Rickrolling, we set the loss weight to and fine-tune the encoder for 100 epochs with a clean batch size of 64. In Villan Diffusion, we fine-tune the model on CelebA-HQ-Dialog dataset [26] with LoRA [23] rank as 4 and the training batch size as 1.
@@
Evaluation Settings. Detection: We select 3000 backdoor samples which contains eight types of triggers and 3000 benign samples for training. The test dataset consists of 3000 backdoor samples which contains other eight types of triggers and 3000 other benign samples. Besides, the target content of the triggers are also different between the training and test datasets. We compare each backdoor detection method in terms of precision, recall, and F1 score, respectively. We also report the inference time for each method. Localization: We test each backdoor attack method with eight different triggers. We conduct experiments to analyze the impact of different similarity thresholds and different similarity calculation tools (i.e., CLIP [37] and DinoV2 [36]). A total of 1000 backdoor samples and 1000 benign samples are used. F1 score is employed as the evaluation metric . The trigger length of Algorithm 1 is set to 1. Mitigation: We test each backdoor method with eight different triggers, and each trigger includes 100 backdoor samples. We utilize two state-of-the-art concept editing methods for mitigating backdoors. The Attack Success Rate (ASR) is computed on each trigger, which is the proportion of the prompt containing the trigger that generates the content specified by the attacker. Besides, we design a novel metric called Average Similarity to Benign (ASB), where we utilize CLIP [37] to compute the image similarity between the benign outputs and the mitigated outputs. We consider ASB is a more strict metric to evaluate the effectiveness of the backdoor mitigation on T2I diffusion models.
@@ 4.2 Detection Results
Results. Tab. 1 and Tab. 2 show the detection performances of F Norm Threshold Truncation (FTT) and Covariance Discriminative Analysis (CDA) under two attack scenarios, respectively. We find that backdoor samples from Villan Diffusion [8] are more deceptive than samples from Rickrolling [43]. This is reflected in the lower F1 score obtained by both detection methods when applied to samples from Villan Diffusion. As shown in Fig. 1, triggers from Rickrolling simultaneously assimilates the structure and intensity of attention maps from other tokens. In contrast, although triggers from Villan Diffusion also exhibit an "Assimilation Phenomenon", each token has variations in the response intensity. This leads to a closer resemblance to the response patterns of benign samples. Besides, CDA shows more robust detection performance on two backdoor attack scenarios, where it all achieves an average F1 score of 88.9 detection results compared to an average F1 score of 86.5 by FTT. In particular, for detecting hard samples from Villan Diffusion, it shows an improvement of 8.9 F1 score compared to the FTT method. Besides, we record the average inference time of the FTT and CDA for detecting a sample on RTX 4090 32GB GPU. As can be seen, both of them perform in real-time, i.e., 9.4 ms for FTT and 11.7 ms for CDA, achieving low computational cost.